Report of the Endowment Fund Task Force (EFTF)

November 30, 1998

Contents

Section  Page
Summary and Preamble  1
Endowments Grants Committee  2
Proposals  2
What gets funded - or not  4
No Appeals  5
Reporting  5
Timeframe  6
Local administration of grants  7
Impact on the Executive Office  7
Communicating with applicants  8
Imagining the EGC at work  9
Suggested form of Summary CV  11
Mandate of this Task Force  12

Summary and Preamble

The Task Force recommends the creation of a CMS Endowment Grants Committee (EGC) as a standing committee of the CMS to administer the distribution of some portion^{1} of the endowment fund income on an annual basis. In general, we support the funding of proposals which contribute to the broader good of the mathematical community as defined in the CMS purpose statements, especially if CMS funds can be leveraged, and we favour funding many smaller projects rather than few bigger projects.

If the CMS does adopt this model or some variant, we recommend that a review of the EGC and its mandate take place after three years. We further recommend that the Board delegate to the EGC committee the authority to make final decisions on funding and provide an annual report to the Board.

1The Finance Committee will be recommending an exact figure.

Endowment Grants Committee (EGC)

The EFTF recommends the creation of the CMS Endowment Grants Committee to administer the disbursement of funds from the income of the Endowment Fund.

The membership of the CMS EGC is to be as follows:

• A current or recently serving member of the Executive.

• A current or recently serving member of the Research Committee.

• A current or recently serving member of the Education Committee.

• A current or recently serving member of the Publishing Committee.

• A current or recently serving member of the Board.

No committee member may be party to a proposal for funding. Members are expected to serve for two years. The EGC is under no obligation to spend all the funds at its disposal. We recommend that funds unspent by the EGC in a given year be reinvested in the Endowment Fund.

Membership of the committee would be suggested for election by the Nominating Committee. The Nominating Committee would be responsible to ensure that the EGC has appropriate representation by language, gender and region.

Proposals

• Proposals must address the goal of the CMS and the CMS purpose statements which we add here.

The goal of the Canadian Mathematical Society is to support the promotion and advancement of the discovery, learning, and application of mathematics.

The CMS statements of purpose are

1. To unify and support Canadian mathematicians through effective communication, broad membership, sponsorship of diverse activities, and partnerships with like professional societies.

2. To support mathematics research through the communication of current research to both the specialist and non-specialist, public recognition of research accomplishments and collaboration with the research institutes and granting agencies.

3. To support the advancement of mathematics education through joint projects with mathematics educators at all levels, promotion of educational advancements, and partnerships with provincial ministries of education and organizations supporting mathematics education.

4. To champion mathematics through initiatives that explain, promote and increase the general understanding of mathematics, provide extra-curricula opportunities for students, and encourage partnerships with corporate, government and not-for-profit agencies.

• Proposals may be fully funded, partially funded, or unfunded. An applicant may be involved in only one proposal per competition as a principal applicant.

• Proposals must come from CMS members, or, if joint, at least one principal applicant must be a CMS member.

• Proposals are expected to be short, two pages at most, in 12 point type, with 1” margins, double spaced. Each proposal must include a summary of 100 words or less, and should indicate to which organization the cheque should be sent in the event the proposal is successful. Each principal applicant must provide a short two-page summary CV. We append a draft of such a summary CV at the end of this report. The CV must include a description of the funding applied for or held by the proposer(s) that has relevance to the proposal. Each principal applicant is required to answer the following question: Have you received funding from the EGC as part of a previous proposal? If the answer to this question is yes, the applicant is required to submit a copy of the report on the most recently funded proposal. [The intent here is to provide an administratively simple checking mechanism.] Failure to include a copy of such a report will disqualify a proposal.

• Four copies of each proposal must be provided by the applicants. This supposes that the EGC opts to use the system of primary and secondary readers suggested by this report in the section entitled “Imagining the EGC at work”. In this case, one copy is for the Executive Office, one is for the Chair, and one each for the primary and secondary readers. Each proposal should contain a stamped, self-addressed envelope which will be used by the CMS to acknowledge receipt of the proposal.

In summary, each submission should contain a self-addressed stamped envelope and four copies of

1. a two-page description of the proposal, including a summary of 100 words or less.

2. A summary CV for each principal applicant, at most two-pages long. Applicants must append a copy of the report on the most recently funded proposal in which they were involved.

• In principle, the committee will favour proposals where CMS funds can be leveraged or where proposals have no other natural funding body to which to apply. Opportunities for leveraging CMS funds and/or alternate sources of funding should be made explicit in the proposal.

• The committee will consider funding proposals for a maximum of three years. However, multi-year proposals must be funded from the funds available to the committee in the year of application.

• The committee will normally consider funding proposals to a maximum of $5K per year and, in exceptional circumstances, up to $10K per year.

• In some cases, it may be that the proposal will generate something of lasting financial value. The committee would find it helpful if the proposers would indicate if this may be the case for their proposal, and declare their intent with respect to that value.

What we may versus what we will not fund.

The Task Force discussed the possibility of forbidding the EGC from funding certain types of potential expenditures and encouraging the funding of other types of expenditure. The discussion was difficult. In general, we believe the EGC ought to be trusted. However, we did agree as follows.

What we won’t fund

The committee will not fund sabbatical salary top-ups. The committee will not normally fund activities that are funded by NSERC or the Institutes.

What we may fund

Course buy-outs and travel costs may be permissible in some circumstances as part of a larger project. However, if a proposal poses a course buy-out(s) explicit written permission for this from the appropriate administrator should accompany the proposal. Subsidizing undergraduate mathematics conferences is permissible, for example to permit the conference to fund some student travel. Funding support for mathematics camps or the like is permissible. Some support for conferences and visitors is permissible. Summer salaries for undergraduates are permissible as part of a larger proposal. However, NSERC now funds summer research students at all insitutions, so we would not expect the EGC to fund this form of activity.

The Task Force does not favour funding regular special sessions at CMS meetings through the EGC, but it may be appropriate to fund extended sessions or satellite meetings held in connection with a CMS meeting via the EGC.

No opportunity for appeal.

The Task Force considered the possibility of an appeals process, but decided against this. There were a number of grounds. We are concerned with our ability to construct a process that would be seen to be fair. We are concerned with the limited ability of the CMS to administer such an appeals process, especially at this early date when we do not know how many proposals we can expect in any given year. We further believe that an appeals process may be unnecessary when there is an opportunity to reapply in a year’s time. On the other hand, we do want the community and the applicants to be encouraged to write the members of the EGC, the Executive and/or the Board with their comments. It may or may not be possible for the EGC to provide feedback to individuals on the structure and/or suitability of their proposals. This seems to depend upon the number of applications received.

We recommend this issue be revisited upon the review of the EGC.

Reporting

• The EGC will publish data on each competition on Camel including the number of proposals and the number of proposals funded. The committee should be aware that their data will be scrutinized to determine the quality of the proposals funded, representation by region, language, gender and whether the funds are spent for research or education.

• Successful multiyear proposals are required to submit one page progress reports, satisfactory to the EGC, before new funds are released.

• All successful proposals are required to submit a one page report at the end of the funding period. No proposals will be funded if they involve individuals who have not reported on prior awards.

• Published materials resulting from a funded proposal, such as advertisements or journal articles, should acknowledge support from the CMS.

• All successful proposals and reports are to be posted on Camel.

Timeframe if the Winter Meeting is used

The Task Force suggests that the EGC meet once a year. On balance, we favour the Winter meeting. If the EGC met at the Summer meeting, proposals involving summer students would have to be made about a year in advance. Proposals scheduled to begin in September might benefit from receiving funding in January, especially if the intent is to leverage funds, and probably wouldn’t suffer from the advance thinking required. It may help those with limited term appointments to achieve further appointment if funding is received in January.

• Call for proposals June 1 with submission date of September 15. Proposals must arrive in the Executive Office before the end of the working day on September 15th or before the end of the last working day prior to September 15th if the 15th falls on a Saturday or Sunday. Late or incomplete proposals will not be considered. At this time, we cannot allow for proposals to be submitted by fax or electronically. We recommend that the CMS allow for electronic submission of proposals as soon as is practical.

• Proposals are sent to the Chair of the Committee by October 1.

• The Chair will forwards the proposals to the primary and secondary readers by October 15.

• Funding decisions to be made at the winter meetings and approved by the EGC, and reported to the Board. The timing here would require the committee to meet the day before the Board.

• The Chair is responsible for conveying funding decisions to the proposers by January 1.

The EFTF expects the EGC to examine its operations and procedures on an ongoing basis and recommend and record such changes to these as are necessary.

Local administration of CMS grants

CMS funding is a grant. CMS funding is not to be used to fund overhead costs under any circumstances. Our expectation is that funds that are not used within the granting period will be returned to the CMS within three months of the end of that granting period. Our expectation is that a local Office of Research Services (or equivalent) will administer grants provided by the CMS for research purposes. Note that the Office of Research Services at many universities wants a copy of proposals either prior to submission or as they are submitted. Our expectation is that grants provided by the CMS for non-research purposes will be administered by the local department, through the Head or Chair, or the equivalent, for example, the director of a school board. Our expectation, in both cases, is that the Head, Chair, or equivalent, will countersign expenditures as they are charged to the project.

Impact of the EGC on the Executive Office

The administration of the EGC will have an impact on the Executive Office (EO) in Ottawa. The EO will be responsible for storing one copy of each proposal received, and forwarding the remaining copies to the Chair. Each proposal will need to be date-stamped. Each proposal will be acknowledged using the self-addressed stamped envelope included with all complete proposals: the acknowledgment will detail whether or not the application was received on time and is judged complete and whether it will be considered, or not. The title, name and address of the principal applicant, and funding requested will be recorded on a summary sheet that will be passed on to the Chair. The Executive Director will be responsible for forwarding funds to successful applicants on the advice of the Chair of the EGC. The EO will be responsible for placing copies of the successful proposals and subsequent reports on Camel, and correspondence with the applicants will be routed through the EO.

The Task Force recommends that funding of up to $1,500 be allocated by the CMS for use by the EO to administer the work of the EGC.

If the EGC uses the system of primary and secondary readers suggested here, then the summary sheet should include columns where the Chair of the EGC can record the primary and secondary reader of each proposal, whether or not the proposal was funded and at what level.

Communicating with applicants

We recommend a postcard be written that will allow personnel in the EO to check off as follows:

• Your application for funding from the CMS has been received, and appears to be complete, thank you.

• We are sorry but we are unable to consider your application for CMS funding because

– your proposal was received in our office after the due date.

– You failed to provide four copies.

– Your proposal violated our requirement that proposals be typed on 8.5x11 paper, using 12 point type, double spaced with 1” margins.

– Your proposal was longer than 2 pages.

– The CV(s) you provided were longer than 2 pages.

Here are suggested wordings for the letters of congratulation and rejection. Rejection.

This year the CMS received x applications for funding various projects in mathematics from around the country. The Endowments Grants Committee, of which I am the Chair, is delegated the authority to adjudicate the funding decisions. Many of the proposals were very good, far more than we had funds to distribute, and we were able to fund only y of them. Unfortunately, your proposal was among those that we were unable to fund. More details regarding the competition will be announced at www.camel.math.ca/ EGC.

Congratulations.

This year the CMS received x applications for funding various projects in mathematics from around the country. The Endowments Grants Committee, of which I am the Chair, is delegated the authority to adjudicate the funding decisions. Many of the proposals were very good, far more than we had funds to distribute. I am happy to be writing you to tell you that your proposal was one of y that we were able to fund.

Enclosed please find a cheque made out to your department for $z. I attach a document that details our expectations for the way in which the funds will be administered. We require you to submit a one-page report by September 15. We also require you to acknowledge the CMS contribution to your project in a visible manner.

The enclosure is the paragraph on “local administration of CMS grants”.

Imagining the EGC at work.

Here are our suggestions for the internal process to be followed by the committee. Proposals are to be farmed out to primary and secondary readers for preliminary ranking and funding suggestions. So each member is to rank and make suggestions for funding one fifth of the proposals as primary reader and one fifth as secondary reader. The EGC is to meet once each year at a CMS meeting to adjudicate proposals.

Let us suppose that our initial call for proposals generates 100 proposals. Each member of the Committee would receive 40 proposals to read, 20 as primary reader and 20 as secondary reader. The Chair would be responsible to record which Committee member is primary reader and which Committee member is secondary reader for each proposal on the summary sheet Each member of the Committee then ranks each of the 20 proposals they receive as primary reader and a recommended level of funding for each one. Each member then does the same for the proposals for which they are secondary reader.

At the meeting, the EGC may decide to proceed proposal by proposal, or member by member. In either procedure, the proposals are divided into three piles:

• If both primary and secondary reader agree that a proposal should go unfunded, that proposal is discarded with no further discussion;

• If the primary and secondary reader disagree on funding a proposal (at any level), that proposal is retained with similar proposals for potential further discussion;

• If the primary and secondary reader agree that a proposal should be funded at some level, that proposal is retained with similar proposals for further discussion.

The Committee then discusses the proposals in the third pile with a view to fully or partially funding them, keeping a running total of funding allocated. In the discussion, it is expected that the primary and secondary reader will explain their reasons for their rankings and recommended funding. It seems clear that the Committee ought not to exhaust the funds available to it without discussing each of the proposals in this pile, unless some method is found of using the rankings to rate the proposals. For example, the committee could decide to fund all proposals where both readers rated the proposal in their top five or so …. If the Committee does not exhaust the funding available on this third pile, it can then review each application in the second pile, in more or less the same way as the third.

Suggested form of CV, draft version

Name: Title of Proposal: Starting Date: Termination Date: Preferred Form of Address (Dr, Mr, Mrs, Ms): Position: Employer: Address:

Telephone: Fax: Email: Web Home Page address: Home Address: Home Telephone: Have you ever received funding from the Endowment Grants Committee of the CMS before? If so, please attach a copy of the report to this form. Please indicate all sources of funding you have which are relevant to this proposal:

CMS Endowment Fund Task Force Rationale, Membership, Mandate and Timetable

Rationale

The CMS endowment fund is at approximately $1.35 million. A commitment has been made to start using the income generated by the endowment once the principal reaches $1.5 million. It is therefore now necessary to develop both guidelines for the use of the income and a strategy to ensure an ongoing revenue stream from the income. The first of these tasks will be done by the ad hoc Committee on the CMS Endowment Fund, and the second by the Finance Committee.

Membership

1. One member from the executive as Chair: Eddy Campbell.

2. Two members from the board: Jennifer Hyndman and Georg Schmidt.

3. Two members from the CMS committee chairs: Bob Rosebrugh and Tom Salisbury.

4. Cms Executive Director as support person: Graham Wright.

Members will be chosen to ensure geographical representation.

Mandate

After broad consultation within the CMS community, to bring to the executive a recommendation on the appropriate administrative structure for the disbursement of funds from the income of the endowment. The Task Force should address in particular:

• Make-up of the governing body and how its members will be determined.

• Terms of reference of the governing body.

• Terms governing the selection of projects and disbursement of funds.

In preparing their recommendations the Task Force should address the following questions:

• What is the process for applying for funding and how will funds be distributed?

• Could awards be made for multiple years (that is, $x per annum for three years)?

• Who will be eligible to apply?

• How should applications be made? What form should they be in?

• What is the time-frame?

• What restrictions might there be on the level of funds requested?

• What role will leverage play (will requests with matching finds from elsewhere be favoured)?

• How will one project be selected over another?

• What kinds of projects will be funded?

• How might this decision fit within the CMS strategic plan?

• Could the funds be used for operating expenses?

Time Frame

• December 97 - executive to approve terms of reference and establish committee

• February 98 - committee membership to be decided

• April 98 - committee to meet in Toronto

• June 98 - committee to consult with Board, Advisory Group and General meeting

• October 98 - committee to present proposal to executive

• December 98 - executive to take proposal to Board